### v1

Abstract
NL2VIS (natural language to visualization) is a promising and recent research area that involves interpreting natural language queries and translating them into visualizations that accurately represent the underlying data. As we navigate the era of big data, NL2VIS holds considerable application potential since it greatly facilitates data exploration by non-expert users. Following the increasingly widespread usage of generative AI in NL2VIS applications, in this paper we present V-RECS, the first LLM-based Visual Recommender augmented with explanations (E), captioning (C), and suggestions (S) for further data exploration. V-RECS’ visualization narratives facilitate both response verification and data exploration by non-expert users. Furthermore, our proposed solution mitigates computational, controllability, and cost issues associated with using powerful LLMs by leveraging a methodology to effectively fine-tune small models, like LLama-2-7B. To generate insightful visualization narratives, we use Chain-of-Thoughts (CoT), a prompt engineering technique to help LLM identify and generate the logical steps to produce a correct answer. Since CoT is reported to perform poorly with small LLMs, we adopted a strategy in which a large LLM (GPT-4), acting as a Teacher, generates CoT-based instructions to fine-tune a small model, Llama-2-7B, which plays the role of a Student. Extensive experiments - based on a framework for the quantitative evaluation of AI-based visualizations and on manual assessment by a group of participants - show that V-RECS achieves performance scores comparable to GPT-4, at a much lower cost. The efficacy of the V-RECS teacher-student paradigm is also demonstrated by the fact that the un-tuned Llama-2-7B fails to perform the task in the vast majority of test cases. We release V-RECS for the visualization community to assist visualization designers throughout the entire visualization generation process.
keywords:
Visualization recommendation, LLM, Machine Learning for Visualization, Chain of Thoughts, Low-cost Generative AI
[…]
During the training phase, a large LLM, GPT-4, acting as a Teacher (B), receives as input triples (D,Q,V) extracted from an available dataset, where D is tabular data, Q is a natural language query, and V is the VegaZero specification for a chart matching the user’s query Q. The model is asked to perform three Chain of Thought (CoT) tasks (A): T1, to explain the reasoning steps that justify the answer V given the query Q; T2, to generate a caption that describes the chart, and T3, to suggest additional useful queries. The responses of the model are combined to create a visualization narrative, VN (C,D). Next, a small model, Llama-2-7B (the Student), is fine-tuned (E) with the enriched dataset of quadruples (D,Q,C,VN). At inference time, the resulting specialized model, named V-RECS, receives as input a pair (D,Q) and generates, along with V, a narrative made of an explanation E, a caption C, and a suggestion S of additional queries. Adopting a teacher-student metaphor, the esteemed professor GPT-4, leveraging its trillion-parameters knowledge of data visualization and natural language understanding, generates examples of visualization narratives for the young student Llama-2-7B. Llama uses this supplementary material to diligently learn the task of producing similar narratives alongside the recommended visualization, and finally gets its doctorate in Visual Recommendation with Narratives, becoming Dr. V-RECS. During the exercise of its profession, will Dr. V-RECS be able to match Professor GPT-4’s abilities? Certainly, its fees are much cheaper.
Introduction
1 Introduction
AI and Big Data development has reshaped how companies make decisions in recent years. A more data-driven approach has become increasingly crucial for companies to make more informed decisions, from big corporations to small and medium enterprises [49]. From marketing [40] to business intelligence, the data-centric strategy represents a terrific opportunity to keep up with a specific field’s evolution and discover patterns that otherwise would be hidden. Notwithstanding, this introduces new challenges and complexity. The data analysis and transformation into actionable insights demand specific domain experts and economic resources. Often, especially for SMEs, these factors represent challenging aspects due to the need for more financial resources to recruit these specialized profiles. Furthermore, the supply of these specialized data scientists often falls short of meeting the burgeoning needs of the industry [^1] https://shorturl.at/ePUXY.
Visualization Recommendation Systems (VRSs) try to fill the gap between non-expert users and data analysis [36], by providing automatic tools to support visual discovery from data across different sectors and applications. Many contributions have been developed in this field (among others, [15, 26, 38, 29]), some of which gained widespread adoption in well-known tools such as Tableau222www.tableau.com, which integrates a functionality named Suggest Me [30]. To date, state-of-the-art VRSs do not achieve full automation, relying on built-in knowledge (such as perceptual rules) or user input to guide the system to generate visualizations, for example by entering natural language queries. This latter category of VRS is referred to in the literature as NL2VIS (Natural Language to Visualization). Many NL2VIS solutions rely on classic NLP (Natural Language Processing) or rule-based techniques [42, 13, 32] to translate the user’s query, called Natural Language Query (NLQ), into insightful visualizations. However, reliance on these techniques poses a significant limitation, as it constrains users to adhere to specific formats (such as the need to clearly specify in the NLQ the type of chart and the variables to be plotted on the axes), thereby restricting the flexibility of expression in their queries. Furthermore, these approaches struggle to manage ambiguous texts, failing to produce a consistent and insightful visualization in many cases.
Following the disruptive use of generative artificial intelligence across application domains, LLM4VIS (Large Language Model for Visualizations) [53] is becoming one of the most promising research areas to advance existing NL2VIS approaches. LLM4VIS aims to use cutting-edge LLM technology to query data sources using only natural language queries and translate them into visualizations that accurately represent the underlying data. Although LLM like Llama [51] or GPT [34], have shown a more advanced semantic understanding of free text [7], they suffer from the lack of interpretability of the model’s responses. Furthermore ”Large" LLM, those involving computationally intensive and expensive architectures containing billions, and recently trillions333as postulated by several sources, among which https://the-decoder.com/gpt-4-has-a-trillion-parameters/ of parameters, also have other contraindications, such as hallucinations [16], the lack of controllability [47] and interpretability [54], and the high computational cost444in https://shorturl.at/puz24 -among other published experiments - it is shown that Llama 2 obtains similar performance to GPT-4 for a fraction of the cost on a number of tasks, hindering their large-scale adoption in applications. In particular, controllability has a notable, and often neglected, impact on applications: if a certain model is based on the adaptation of a proprietary LLM, such as GPT, the evolution of the original LLM [^5] For example, there are notable differences between GPT-3 and GTP-4. could significantly impact the adapted model. This means that developers of a model based on a proprietary LLM do not have full control of their system.
Finally, although LLMs have proven great capabilities on tasks like text, images, or code generation, it has been observed that they are much less proficient on other tasks, such as data visualization and analysis. In this regard, Podo et al. [35] observed how, even though GPT-3.5-turbo can already perform data visualization tasks, it still produces a high rate of errors and poses a problem of response verification, especially when the user is not sufficiently expert to evaluate the result. Unlike text and images, which can be easily verified by any user, or code, which can be verified by compiling and then validating the expected output, evaluating the quality of a recommended visualization is more challenging. If users do not know what to expect, or they are not domain experts, they may perceive the model’s output as accurate, even when incorrect.
To mitigate some of the outlined problems, in this paper we propose V-RECS, a LLM-based Visualization Recommender enhanced with Explanations (E), Captioning (C), and Suggestions for further data exploration (S). The data visualization narratives E and C mitigate the problem of response verification, while S support non-expert users to further explore the available data. The generated model, based on a fine-tuned ”small" LLM, Llama-2-7B, also mitigates cost and controllability issues [^6] Llama-2-7B has seven billion parameters, while, e.g., GPT-4 has one trillion. Furthermore, Llama-2-7B is available free of charge for research and commercial use; in contrast, GPT-n models are proprietary, and their inner workings are hidden from the public https://shorturl.at/bdoOZ.
V-RECS’ training methodology is based on Chain of Thoughts (CoT), a strategy to instruct a large LLM to break down its responses into reasoning steps[55]. Since CoT has been shown not to work well with small LLM [53], we adopt a teacher-student paradigm in which GPT-4, the teacher, is asked to enrich with visualization narratives an existing dataset of users’ queries and corresponding charts, leveraging the CoT strategy. Next, the enriched dataset is used to fine-tune a small open-source LLM, Llama-2-7B, to perform visualization recommendations along with textual narratives to explain and further explore the input data. V-RECS exploits the synergies between LLM and visualization to improve the generation of insightful visual representations from data, while ensuring that the user has full control over the reliability of what is suggested by the Visual Recommender. The main contributions of this work are:
  • V-RECS, the first LLM fine-tuned for the visualization field that performs visualization recommendation along with the production of insightful visualization narratives;
  • Low-cost and controllable solution: V-RECS is based on fine-tuning the open source Llama-2-7B, which was estimated to be 100 times less expensive than GPT-4, along with the other previously summarized advantages of low-parameter Open-AI models 777amongst the many studies, see e.g.,https://tinyurl.com/56kpjd7r. Despite leveraging a small LLM, V-RECS achieves performances comparable to GPT-4, as shown in Section 6.
  • LLM4VIS methodology: we present an end-to-end solution for LLM4VIS, contributing a novel approach that bridges the gap between generative artificial intelligence, explainable machine learning, and the visualization field.
  • A comprehensive comparative quantitative and qualitative evaluation experiment, to study how the proposed V-RECS model and GPT-4 perform the visualization task.
We release V-RECS for the visualization community on a GitHub repository [^8] https://github.com/lucapodo/V-RECS.git.
The paper is organized as follows: in Section 2 we summarize the main concepts of LLM. Section 3 is dedicated to the state-of-the-art in NL2VIS and LLM4VIS. Section 4 presents a formal description of the LLM4VIS problem. Section 5 describes the proposed V-RECS model and the LLM-based methodology to develop it. Section 6 reports the results of quantitative and qualitative evaluation activities. Finally, we present a discussion and our concluding remarks in Section 7.
2 Background
LLMs are powerful machine learning models based on the concept of pre-training followed by fine-tuning. During pre-training, the model learns representations of language through unsupervised learning on extremely vast corpora of textual data. This phase is crucial for the acquisition of general linguistic knowledge. Pre-training is based on Transformer-based architectures [23], capable of capturing long-range dependencies and contexts, essential for understanding the complexities of human language. Following pre-training, fine-tuning tailors the LLM to specific downstream tasks by exposing it to labeled data and adjusting its parameters through supervised learning. This process adapts the model’s learned representations to the specific task at hand, enhancing its performance. Recently, LLMs have been adapted to process and generate different types of unstructured data beyond text, such as images and videos, music, climate data, code, and visualizations, which is the main focus of this paper.
Another relevant concept in LLMs is Prompt engineering [39], a technique involving the design of effective instructions (prompts) or input examples to guide the behavior of language models during fine-tuning or inference. It aims to design instructions for the LLM that elicit the desired responses from the model while minimizing ambiguity and bias. Prompt engineering is a key area of research to extend the capabilities of state-of-the-art LLM. We can distinguish between two main prompting techniques: zero and few-shot prompting. Zero-shot [20] represents a paradigm where the model is tasked with performing a given objective solely based on a task description without explicit examples illustrating how to accomplish it. This methodology leverages the model’s embedded knowledge to generate appropriate responses. Conversely, few-shot learning [37] enhances the LLM capabilities by augmenting the task description with a small set of example inputs paired with their corresponding outputs.
3 Related work
A Visualization Recommendation System (VRS) is a system designed to suggest insightful visualizations given a dataset. As outlined by Podo et al. [36], there are two classes of VRS: task-agnostic and task-aware. The task-agnostics [15, 11] require solely the dataset as input without additional guidance to generate insightful visualizations. Most of the approaches rely on classical machine or deep learning methods. Recent advancements have explored leveraging LLMs [10, 53] to perform this task. In this case, the visualization recommendation process is split into multiple steps (e.g., dataset filtering, dataset description, column selection), with the LLM involved in all or some of these steps.
While attaining full automation in visualization recommendation pipelines would signify a remarkable milestone, it is still in the early stages, facing several challenges that have not yet been explored [36]. Conversely, task-aware VRS requires user input or guidance, typically as user queries, to generate visualizations. Task-aware VRSs are extensively studied in the literature and fall within the domain of Natural Language to Visualization (NL2VIS). In NL2VIS, the system is guided by a natural language query that tells the model what to explore in a dataset (e.g.: What are the payment date of the payment with the amount paid higher than 300 or with payment type is check, and count them by a bar chart, order by the y axis from low to high?). Recently, the field has been significantly influenced by LLMs, which demonstrate considerable ability in understanding human language, allowing them to address the challenge of ambiguity inherent in user expressions that currently limit existing approaches.
3.1 NL2VIS approaches
The development of NL2VIS has witnessed significant advancements, primarily through classical NL techniques.
As discussed by Setlur et al. [42], the authors introduce an interactive visualization tool that employs a hybrid approach integrating NLP and decision rules. This system combines a probabilistic grammar approach with predefined rules that can be dynamically updated based on the underlying data. Another example is proposed by Srinivasan et al. [45], where the authors present BOLT, a system based on traditional NLP techniques to map user queries to prevalent dashboard objectives and generate appropriate visualizations. In contrast, the solution proposed by Narechania et al. [33] focuses on generating a single visualization rather than suggesting a complete dashboard.
While these approaches provide reliable solutions, they impose input constraints by asking users to adhere to specific input templates rather than freely using natural language (for example, explicitly naming the variables to be considered in the chart, avoiding semantically ambiguous sentences, etc.). A departure from these syntactic approaches is NcNet, proposed by  [28]. The authors propose to use a seq2seq neural network model to translate user queries into visualizations, leveraging a novel visualization grammar named VegaZero. VegaZero streamlines the complex Vega-Lite specification into a more intuitive sequence-based keyword format. Generating a plain text sequence proves to be more manageable for a seq2seq model, in contrast to the challenges associated with hierarchical outputs, such as for Vega-Lite.
Despite the advancements introduced, classic NLP systems encounter various challenges. Primarily, they depend on the well-structured syntax of user queries for effective processing. Furthermore, failure to understand the semantics of the query fails to understand what the user wants to explore.
3.2 Large Language Models for visualization
Within the field of NL2VIS, Large Language Models for Visualization (LLM4Vis) is emerging as the new frontier for translating user queries into consistent visualizations. LLM4VIS seeks to integrate LLMs into data visualization pipelines to achieve full support for visualization automation. They have shown an increasing proficiency in comprehending the underlying semantics of human instructions [14], representing a compelling opportunity to overcome the limitations inherited by the ”classic" approaches illustrated in the previous section. While this technology is still in its early stages and requires further exploration and study, it nonetheless presents an emerging opportunity for the visualization field, as discussed by Schetinger et al. [41]. We can categorize the LLM4VIS approaches into two main classes. The first class encompasses all the methods that involve the model in its standard configuration (the so-called ”plain foundation models") and use prompt engineering techniques to improve the model’s performance. Conversely, the second class includes all the approaches that involve a "fine-tuning" phase to specialize the foundation model toward a specific set of tasks.
LLM4VIS prompt tuning approaches A novel method for translating natural language queries into visualization is proposed by Chen et al. [4]. The authors involve Codex - a pre-trained LLM developed by OpenAI for natural language to code translation  [^9] https://openai.com/blog/openai-codex - configured within a few-shot learning prompt setup. They propose a two-step strategy to recommend a visualization. Firstly, the input is shaped as examples of natural language to SQL pairs along with the user’s query, letting the model generate the most appropriate SQL. These examples serve as reference points, guiding the model in addressing new user queries effectively. Then, the response generated by the model is translated into Vega-Lite specifications using a rule-based approach [5]. Maddigan et al. [31] present a comparable study involving Codex, GPT-3, and ChatGPT, assessing zero-shot prompting to evaluate the quality of these models on generating visualizations as Python code. A notable limitation of these works is the lack of comprehensive discussion regarding results. Instead, they rely primarily on a small number of qualitative evaluation experiments, merely providing an overview of potential implementations. Li et al. [22] introduce Sheetcopilot, an LLM-based system that uses natural language queries to drive user interaction within spreadsheet software, such as table filtering or chart generation. Sheetcopilot involves in-context learning prompting technique to interact with all spreadsheet functions, employing GPT as model. It utilizes the user’s query alongside a prompt encompassing the model task, a comprehensive list of software functions (i.e., atomic actions), and an output format. With these inputs, Sheetcopilot translates the user request into a systematic plan, outlining step-by-step actions to fulfill the user’s query effectively. A similar approach is Datacopilot introduced by Lin et al.  [24]. Datacopilot is a method to generate automated interfaces and autonomous workflows for human-table interaction, solving different data-related activities, such as acquisition, processing, forecasting, table manipulation, and visualization. It involves GPT in in-context learning to generate and deploy a data processing workflow using small input guidance. However, both Sheetcopilot and Datacopilot, while capable of generating visualization, are not optimized for this task. Their primary focus lies in generating instructions or commands to fulfill user requests within software environments like spreadsheets or toolsets rather than optimizing for insightful visualization.
Overall, while existing prompt tuning approaches offer cost-effective deployment at inference time, they come with limitations. Relying solely on prompt engineering results in a lack of control over the model’s behavior in task execution. Additionally, the majority of these methods depend on proprietary models like GPT. Any updates made to these models may introduce new behaviors, potentially undermining the efficacy of the proposed methodologies.
LLM4VIS fine tuning approaches. A distinct approach to aligning a Large Language Model (LLM) with the visualization task involves fine-tuning the model. Zha et al.  [56] introduce TableGPT, a fine-tuned model designed for querying data tables solely through natural language queries. To mitigate input token length constraints, the authors employ an encoder to preprocess the table, extracting its vector representation. Subsequently, the embedded representation alongside the user query is fed into the LLM, which generates a command set and textual reply. Unlike the previous approach, which focuses on table interactions workflow, Tian et al. [50] propose ChartGPT, an LLM strictly tailored for visualization recommendation. The authors fine-tuned the open-source model FLAN-T5-XL [8], by splitting the data visualization generation problem into two broad tasks (i.e., data transformation and visualization recommendation), each consisting of three sub-tasks. ChartGPT focuses solely on the visualization recommendation task, offering a novel LLM adaptation strategy. However, it lacks a comprehensive visualization narrative. While it can suggest optimal visualizations based on user queries, it fails to explain the reasoning behind the recommendations.
3.3 Generating visualization narratives
Once a model produces a visualization or provides a response to a user query, end users face considerable difficulty in determining the accuracy of the provided output. Compared to textual responses, verifying visualizations poses a more difficult challenge. While tasks such as text summarization, content generation, or code generation allow for relatively straightforward output verification by end-users, the same cannot be said for visualizations because non-expert users may inadvertently consider misleading representations as accurate. Consequently, the production of a ”visualization narrative" (a description of why a recommended visualization fits a user’s intentions implied by the natural language query) emerges as crucial. The task of generating a visualization narrative delivers a textual elucidation of the underlying process behind the visualization generation process performed by the model. This not only enhances transparency but also reduces the potential for misinterpretation or erroneous acceptance of visualized data.
A first approach in this direction has been proposed by Sultanum et al.  [46]. The authors propose Datatales, a method to generate a visualization narrative when given a visualization and an annotation as input. Once the system receives the input, it is passed to GPT-3, which provides a textual narrative based on a simple prompt. Another similar approach is proposed by Ko et al. [19]. The authors present a VL2NL framework to generate a visualization narrative given a VegaLite specification. It accepts as input solely the specification and produces a visualization description and a data description as output. Additionally, it provides a second output, representing a possible query that a generic user could use to generate the visualization. Note that both Datatales and VL2NL perform a different task from NL2VIS, since the visualization is an input, rather than an output.
Unlike all the methods in the literature, we propose V-RECS, a model capable of not only generating a consistent visualization, as proposed by Tian et al. [50], but also explaining the generated output by integrating visualization narrative capabilities, as suggested by Ko et al. [19]. V-RECS aims to combine the strengths of the existing strategies to deliver a system capable of generating and explaining a visualization in a fully-flagged automatic way. Furthermore, V-RECS is designed to be independent of additional software or tools to produce the expected output, as seen in [22, 56, 4]. Instead, we perform all the operations during the reasoning process and provide a verification involving the visualization narrative. V-RECS provides a methodology to adapt smaller models to perform as the larger ones, as proved in the evaluation. Remarkably, V-RECS leverages a ”small" and open-source LLM model, Llama-2-7B, thus mitigating cost, privacy, and controllability issues mentioned in Section 1.
4 Problem formulation
We can formulate the goal of LLM4VIS systems as maximizing the probability of suggesting the best visualization to a user, given a data table $T$ and a user query $Q$:
$\centering f_{i}(Q_{i},\mathfrak{T})\rightarrow argmax_{V_{i}}P(V_{i}|\mathfrak{T},Q_{i})\quad where\quad V_{i}:\{v_{i,j}|0<j\leq n\}\@add@centering$ (1)
where $Q_{i}$ is the user query at time $i$, $\mathfrak{T}$ is the table representation passed to the model and $V_{i}$ is the output visualizations set (a single visualization or a dashboard). Note that, commonly, $V$ is not directly a chart, but rather, the specification of a chart in a visual representation language, typically, a JSON-based language like Vega-Lite and the simpler VegaZero [28].
The previous is a general formulation that also applies to NL2VIS systems. LLM4VIS systems introduce a series of intermediate tasks, illustrated in Figure 1. First, it is crucial to note that the model does not directly receive the original data table $T$ as input, but rather $\mathfrak{T}$, which results from a table transformation operation. Given the LLM’s limited input size, depending on the number of input tokens a model can accept, it becomes essential to pre-process the original table into a format suitable for model input. To face this problem, for example, Zha et al. [56] transform the table using an encoder; differently, Tian et al.  [50] propose to use only the table description with some rows to provide the model a better context of the data. Following the table transformation, the input is shaped by a prompt function $P(u_{i},\mathfrak{T})$ (see Section 2) that maps the query and table representation to a designed prompt template.
Figure 1: V-RECS model task workflow. The three green boxes represent V-RECS’ specific tasks, while the others are common to all LLM4VIS models.
After that, the LLM should perform the following tasks to produce $V$ (as shown in Figure 1, left):
  • columns selection. Once the model receives the prompt as input, the model has to determine the most meaningful and prominent columns from the data table based on the user query.
  • data transformation. After the model has chosen the columns based on the user query and the column’s data type, the most appropriate data transformation functions should be selected. For instance, the model should determine whether to aggregate (e.g., min, average, max) different data columns or values.
  • visualization recommendation. Based on the output of the previous steps, the model must choose the best visualization elements (axes, marks, presentation, etc.) to map the data and generate the visualization as code (e.g., matplot python code to produce a bar chart), image (e.g., bar chart as pixel representation), or grammar (e.g., a visualization expressed using VegaLite specifications).
In addition to the above tasks, in this paper we propose that the LLM model should also perform three additional tasks (see the green boxes in Figure 1):
  • visualization caption. The model should generate descriptive captions for the produced visualizations. These captions elucidate the visual representation’s content and meaning, helping the user understand the data shown.
  • visualization explanation.Subsequently, the model should explain why specific columns have been selected from the available dataset. This transparency is essential for users to comprehend the decisions behind the model’s choices, enabling them to verify the relevance and accuracy of the generated visualizations.
  • queries suggestion. Finally, the model should offer suggestions for additional data queries, guiding users on where to delve deeper into the dataset for further insights.
We name V-RECS the specialized LLM4VIS model able to generate visualization recommendations along with the above tasks, referred to as generation of visualization narratives. Accordingly, the formal specification of the V-RECS model task is as follows:
$\centering f_{i}(Q_{i},\mathfrak{T})\rightarrow argmax_{V_{i}}P(V_{i},N_{i}|\mathfrak{T},Q_{i})\quad where\quad V_{i}:\{v_{i,j}|0<j\leq n\}\@add@centering$ (2)
where $N_{i}$ denotes the rich narrative accompanying the generated visualization.
5 V-RECS Model
In this section, we describe the methodology to implement V-RECS. The proposed methodology is based on fine-tuning a small model in a supervised manner, using an existing dataset enriched with visualization narratives generated by a more powerful model. The methodology is depicted in Figure V-RECS, a Low-Cost LLM4VIS Recommender with Explanations, Captioning and Suggestions. (the teaser). First (box A, B of Figure V-RECS, a Low-Cost LLM4VIS Recommender with Explanations, Captioning and Suggestions.), GPT-4, acting as a Teacher, is instructed to generate visualization narratives leveraging the Chain-of-Toughts technique proposed in [19]. The Teacher receives as input triples from an existing dataset, Ncnet [28], composed by a dataset D, a user’s query Q, and a visualization V matching the user’s needs expressed by Q. Based on the CoT approach, we have defined three tasks, each with a different goal. The first task, T1, aims to explain why a visualization V is a correct answer given the user’s query Q, while the T2 goal involves describing the visualization. Differently, the goal of T3 is to suggest potential areas of interest for further exploration to the user. Next, the teacher-generated responses for the three tasks are combined using a custom prompt template for the subsequent fine-tuning phase (boxes C and D in Figure V-RECS, a Low-Cost LLM4VIS Recommender with Explanations, Captioning and Suggestions.).
Then, the responses produced by the larger model are used as ”expertise" to teach a smaller LLM model (the Student LLama-2-7B) to behave similarly (box E). At inference time (right side of Figure V-RECS, a Low-Cost LLM4VIS Recommender with Explanations, Captioning and Suggestions.), the fine-tuned model, named V-RECS (Visualization Recommender with Explanations, Captions, and Suggestions), receives as input a (D,Q) pair and recommends a visualization V along with a visualization narrative composed by an explanation (E), a caption (C), and suggestions for further explorations (S). To instruct the models we use the zero-shot prompting strategy (see Section 2) for the Teacher and fully-supervised training for the Student. At inference time instead, V-RECS works in zero-shot. In what follows we explain the methodology in more detail.
5.1 Model training methodology
As described in Section 4, the goal of V-RECS is to recommend visualizations, along with rich narratives to explain and extend recommended visualizations. Furthermore, we also aim to build our solution by leveraging a small and open-source LLM model, the advantages of which we have already described in Sections 1 and 3. Smaller models specialized in a specific task, with comparable performance to larger ones, represent a terrific opportunity to reach the same results at a minor cost, while leveraging an open source software. In the literature, approaches aimed at enhancing the capabilities of LLMs can be broadly categorized into those leveraging the model in its existing form, primarily through prompt tuning, such as Chain-of-Thoughts (CoT) [55], and more advanced techniques involving updating the model weights, like the fine-tuning (see Section 3.2). CoT is one of the most popular methodologies among the former approaches. It pushes the model to emulate human-like cognitive processes by telling it to break down complex problems into more manageable sub-problems, enhancing response accuracy. CoT can be classified as zero-shot [20] and few-shot [55]. The few-shot configuration necessitates a set of question-answer pair examples in addition to the new one. In this case, the prompt typically comprises multiple question-reasoning steps-answer triplets, guiding the model through problem-solving. Conversely, the zero-shot configuration requires a specific text formulation, ’Let’s think step by step,’ within the prompt, telling the model to split the problem into sequential sub-tasks. For instance, a query like ’Q: how much is 2+2? Let’s think step by step’ drives the model to break the question and solve each part individually, culminating in a final response. Unlike the few-shot variant, zero-shot CoT does not provide examples of the problem-solving process but only relies on the model’s embedded knowledge. CoT has demonstrated substantial advantages, particularly in inferring new tasks with minimal effort, without necessitating structural changes to the underlying model architecture [55]. However, its application to models with fewer than 100 billion parameters has been associated with logical inconsistencies during the reasoning process [55]. To address these discrepancies, fine-tuning smaller models emerges as a viable solution to enhance the proficiency of small models in solving complex tasks, such as data visualization. Nonetheless, this approach presents its limitations. Fine-tuning smaller models on complex tasks, using question-answer pairs alone, devoid of reasoning steps, often yields incorrect responses. To overcome these limitations and augment the capabilities of smaller models in data visualization tasks, we propose leveraging a novel approach known as Fine-tuning based on CoT [19]. This method combines both CoT and fine-tuning by enriching question-answer pairs without reasoning steps with the sub-tasks needed to solve the problem posed, improving the reliability of a smaller model in producing correct outputs [1]. The process involves exploiting the embedded knowledge of a larger ’Teacher’ model, such as GPT-4, in a zero-shot CoT configuration, to provide responses enriched with reasoning steps. Subsequently, a smaller ’Student’ model is fine-tuned based on these responses, effectively leveraging the Teacher’s knowledge. For example, for the question: ’A juggler can juggle 16 balls. Half of the balls are golf balls, and half of the golf balls are blue. How many blue golf balls are there?’, the Teacher is tasked with a zero-shot CoT, providing logical steps to produce the expected solution such as: ’Step 1: There are 16 balls in total. Step 2: Half of them are golf balls. Step 3: Therefore, there are eight golf balls. Step 5: etc.’. Then, the response generated and the original question are used to fine-tune the Student model, thereby leveraging the Teacher’s knowledge to enhance the final performance of the generated model.
While fine-tuning based on CoT has been used to augment datasets for solving scientific problems [25], its application to visualization tasks remains unexplored. Building upon the methodology by Ko et al. [19], we have devised a custom CoT strategy for the V-RECS problem domain, involving three ’Teachers,’ each tasked with addressing specific challenges related to the V-RECS tasks, including visualization recommendation and captioning, explanation of the process of column selection in T and filtering, and suggestion of additional questions for data exploration. The Teachers’ tasks are delineated in Table 1, with variations based on the input and intended goal.
Finally, the responses generated by the three Teachers, along with the original query Q and the expected visualization V are combined according to a custom prompting strategy used to fine-tune the Student, as detailed next.
Table 1: Description of the Teacher’s tasks
Task Input Goal Description
T1
Query,
VegaZero,
Dataset
Explain
Explain to the user why the features used in
VegaZero are the most appropriate given all
the dataset features and the user’s query
T2 VegaZero Describe Describe the content of the visualization
T3 VegaZero Suggest
Discover the most insightful questions to
suggest to the user for exploring the data
5.2 Implementation details
In this section we provide the relevant implementation details to replicate our methodology. Additional details can be found in the annexed Supplementary Material and at the V-RECS GitHub  [^10] https://github.com/lucapodo/V-RECS.git.
Visualization representation A challenging aspect concerns the choice of a visualization description method. A visualization can be represented as a declarative specification, a bitmap image, or through its code. The first type is based on visualization grammars, like the popular Vega-Lite. The second type considers the visualization as an image, and the third type represents the visualization as a set of instructions, for example, using Matplotlib in Python. Pixel-based visualizations would add limitations and complexity to our approach, as LLMs are primarily trained using textual data, while limiting interactivity and deployability. Similarly, employing a code-based visualization method would hinder the creation of a pure LLM approach, requiring a specific code environment to render and generate visualizations. Differently, a grammar-based visualization can be considered agnostic to the used rendering environment, making it more general than code, while still formatted as structured text. For this reason, we chose grammar-based visualizations as the most suitable for our scenario. Specifically, we have opted for VegaZero, proposed in [28].
Selection of the training dataset Various datasets have been proposed in the literature for Natural Language to Visualization (NL2VIS) tasks, with NvBench [27] emerging as the most prominent and widely adopted. However, it exhibits limitations such as a lack of complexity in the visualization types, overly explicit natural language queries (except for some instances pre-classified as ”hard"), and the absence of support for visualization narratives or reasoning steps. While other datasets like NLVCorpus [44] or Quda [12] provide alternatives, they also come with their own constraints. In particular, NLVCorpus includes only 893 visualization-query pairs, and Quda does not provide the chart associated with the queries. In the absence of a dataset capable of supporting all our requirements, we opted to choose NVBench as the best fit for our purpose. We chose the NvBench version provided by NcNet, as it aligns closely with our task requirements, providing the queries, visualization, and data sources. Specifically, this version excludes visualizations generated from multiple data tables and adds VegaZero representations tailored to each query. Other datasets, such as VisText [48] or VIS30k[3], lack the query-vis-data triples and focus solely on caption generation.
Teacher and Student prompting GPT-4 has been selected as the Teacher and prompted with three distinct tasks (T1, T2, and T3), described at a high-level in Table 1, based on the input and task delivered. Furthermore, leveraging the work of Ko et al. [19], we provide each prompt with in-context reference to the VegaZero template to help the model generate consistent visualization. Each task is tailored for Zero-shot CoT supplemented with strategically embedded sub-tasks, denoted here as ”steps". These steps act as ”guiding principles", offering the Teacher subtle cues on approaching and unraveling the task. The steps vary for each teacher, based on their task.
Figure 2 shows the detailed description of the prompt template for T1 (Explain), split into three parts and related steps. The other prompt templates, along with a more detailed description, can be inspected in the Supplementary Material.
Figure 2: Structure of the T1 prompt. The related task is described in Table 1. It has three parts: A encourages the LLM to reason ”step-by-step" as per the Chain of Thoughts method, B includes the description of sub-tasks (steps) to guide the model, and C is the response template.
Next, all the teachers’ responses $R_{i}$ are combined into a prompt template to instruct the Student, consisting of the header, the input, the body, and the response, as follows:
  1. 1.
    The header includes the explicit task the model is supposed to pursue. It drives the model to understand what it has to achieve. Along with the task, this part also includes the VegaZero template structure and, finally, the "Let’s think step by step" formula [20] which is the basis of the Chain of Thought prompting method.
  2. 2.
    The input includes the user query and the dataset description. The dataset is embedded using only the data feature names and related types.
  3. 3.
    The body consists of the intermediate or reasoning steps the model should follow to select proper visualization properties, such as the aggregation function. This part is crucial to provide additional information to the end-user so that they can understand the model’s design choices and generate the final response. Adding this part enables the explainability of the generated response.
  4. 4.
    The response is the last part of the prompt, where we provide the structure of the answer that must be presented by the model to the end-user. It includes the VegaZero specification V to generate a chart, and the E,C,S triple that makes up the narrative explaining the recommended visualization.
The Student prompt template, along with a filled example of training instance, is depicted in Figure 3. The left-hand side of Figure 3 shows how the different components of the Teachers’ responses contribute to filling the student prompt.
Figure 3: Fine-tuning prompt template (Student prompt). The leftmost box shows how the responses R1, R2, and R3 contribute to filling the different parts of the template. The rightmost box is a (partly) filled template for a specific training/testing instance.
Fine-tuning The last step is fine-tuning the model. As a student, we have used Llama-2-7B [51], the smallest from the Llama2 family. To fine-tune Llama-2-7B, we generated a dataset comprising 2938 triples, each consisting of a dataset D, a user’s query Q, and a VegaZero specification V, following the same distribution of query complexity of the original dataset. Drawing from the QLora technique [9], our training approach went beyond attention modules [52] to incorporate linear components. QLora has shown very promising results during training. With reference to [9], when configuring Lora, we opted for an $r$ value of 64 and an $\alpha$ value of 128. During fine-tuning, we employed a batch size of 4, a learning rate of $1e^{-4}$, and utilized the AdamW optimizer.
The resulting fine-tuned model is V-RECS. The V-RECS model was deployed on AWS Sagemaker with an A10G Nvidia GPU, fed using only the user’s query and the dataset. Differently, GPT-4 operated in a zero-shot CoT, with prompt mirroring used during V-RECS training.
Figure 4 shows an example of V-RECS response at inference time, when given a dataset D and query Q.
6 Evaluation
This section presents a comprehensive evaluation of V-RECS, encompassing quantitative and qualitative assessments. In both types of evaluation, we compare the results generated by V-RECS with the ones generated by GPT-4. While a baseline evaluation was initially planned with Llama-2-7B (the un-tuned model), extensive experimentation with prompt templates and incremental prompt construction yielded unsatisfactory results, leading to its exclusion from further evaluation.
Figure 4: Example of V-RECS response given a dataset and a user’s query
The evaluation was carried out using EvaLLM [35]. Similar to the popular ISO-OSI stack, EvaLLM introduces a layered evaluation framework where each layer’s output serves as the input for the layer above it, including five layers. The lower layers focus on the structural aspects of the visualization, while the upper layers delve into more semantic and insightful properties.
  1. 1.
[…]
    The code layer serves as the foundational level, evaluating fundamental structural properties.
  2. 2.
[…]
    The representation layer addresses core visualization properties based on established representation rules.
  3. 3.
[…]
    The presentation layer assesses the perceptual quality of data presentation.
  4. 4.
[…]
    The informativeness layer gauges the intrinsic quality of visualizations in terms of insights and adherence to best practices in visualization literacy.
  5. 5.
[…]
    Lastly, the LLM layer evaluates the computational and strategic costs of generating specific visualizations.
Furthermore, as detailed by Podo et al. [35], the EvaLLM stack introduces twelve different levels within each layer, each evaluating distinct properties.
Since EvaLLM is a framework, it must be specialized for specific experiments. For the purpose of evaluating V-RECS we chose to focus only on seven of the twelve levels proposed in EvaLLM. Additionally, we introduced a new level, ”Caption quality". We implemented the seven selected levels as follows:
[…]
Syntax correctness: For this level, we employed the Altair Python library’s render to validate VegaLite specifications generated by GPT and a parser developed by VegaZero authors for V-RECS. The evaluation metric used to compare correct specifications against the total evaluated examples was accuracy.
[…]
Note that all the above levels can be evaluated quantitatively, except for Significance, which we assessed with a qualitative evaluation by human judges.
6.1 Quantitative evaluation
For the quantitative evaluation, we used 1800 samples drawn from the original ncNet dataset [28], not used in the training phase, augmented with the generated visualization narratives, as described in Section 5. This experiment aimed to prompt both GPT-4 and V-RECS models with user queries and datasets, subsequently evaluating the responses generated by each model. These responses encompassed the recommended visualizations V and the accompanying narrative, allowing for a comprehensive performance assessment. In detail, the experimental setup involved prompting the models with a user query, dataset, and task description. Conversely, the prompt for GPT-4 was more elaborate, including the task, zero-shot CoT formulation, a set of tasks guiding its response generation, and the desired response format121212i.e., we use a template format similar to those used for the three teachers, see Section 5.2. As shown in Figure 5, the two models exhibited comparable performance across these levels, with GPT-4 showcasing a slight advantage from +3% to +4% accuracy in what is comparable performances between the two models. However, a notable discrepancy emerged in assessing axes quality, where GPT-4 appeared to significantly underperform compared to V-RECS (75% against 93% accuracy). Specifically, GPT-4 frequently exhibited errors such as axis inversions, a phenomenon previously discussed in literature [35]. Despite this, both models demonstrated comparable accuracy across the first six levels, indicating comparable efficacy in generating meaningful visualizations and narratives. Furthermore, it is essential to consider the efficiency aspect of our evaluation. Efficiency, depicted for all dimensions in Figure 5 (right), entails comparing the accuracy achieved by each model at each level, while also considering the model complexity. Notably, our findings suggest that results comparable to GPT-4 can be achieved with a much smaller and more manageable model, which is a relevant advantage in practical deployment scenarios.
Figure 5: Results of the EvaLLM evaluation experiment
6.2 Qualitative evaluation
For the Significance layer, we have designed a qualitative experiment involving 50 randomly selected samples (not used for training) evaluated by 15 human annotators. Importantly, we maintained the original query difficulty distribution (as categorized in the source dataset), ensuring that our evaluation was representative of real-world scenarios. To facilitate the evaluation process, we developed a web application designed specifically for conducting a comparative study  [^13] https://vrecseval.streamlit.app/. This application allowed us to present each volunteer with pairs of responses from the two models for the same query and dataset. The identity of the models’ responses was hidden from participants, ensuring unbiased evaluations. Each participant evaluated the responses using a Likert scale, ranging from 1 (Completely Meaningless) to 5 (Completely Meaningful). This scale provided a structured framework for capturing the nuanced differences in the quality and relevance of the responses. Firstly, each participant was assigned a subset of 20 samples from the total pool of 50. This ensured that each sample received multiple evaluations from different participants, thereby mitigating individual biases. To minimize the impact of any outliers or inconsistencies in the evaluations, we enforced a minimum threshold of three unique evaluations per sample.
Figure 5, in the Significance layer, shows a slight advantage of V-RECS. This is better shown in Figure 6, where results are distributed according to the complexity of the query, and categorized into four levels as in the original dataset. As expected, the quality of answers worsens with the complexity of natural language queries for both systems, with V-RECS performing slightly better except for the extra hard sentences.
Figure 6: Results of the qualitative evaluation by complexity of the natural language query
Upon a subsequent models’ evaluation by the authors, we noticed that the explanations and captions produced by the two systems are slightly different but equally satisfactory, although V-RECS provides a more comprehensive explanation of the reasoning steps. This result mitigates some of the concerns reported in the user study by Kim et al. [18] in terms of lack of depth and critical thinking capabilities while explaining a visualization design. Contrarily, GPT-4 often surpasses V-RECS in suggesting additional questions, presenting greater diversity and serendipity. We believe this difference can be reduced in our future research by improving the fine-tuning prompt strategy for the T3 task. We also noted a couple of hallucinations in GPT-4, while in two cases V-RECS was able to produce a good narrative but, due to some problem in handling data values, could not generate a visualization.
7 Discussion and Future work
This paper has proposed the first LLM-based visual recommender, V-RECS, capable of providing explanations, captioning, and suggestions for further analyses, and has demonstrated its capabilities. Remarkably, V-RECS is based on fine-tuning a ”small", open-source model, Llama-2-7B, while achieving comparable performance to a much larger, proprietary model, GPT-4.
At the same time, in this section, we list some limitations that V-RECS presents to encourage further research on the topic and opportunities that the availability of V-RECS enables for visualization and visual analytics researchers and practitioners.
7.1 Limitations and challenges
Here we report limitations concerning V-RECS and the methodology it is built on. More specifically we report on:
Visualization Literacy Coverage: The fine-tuning of V-RECS and quantitative experimentation are at this moment based on the NCnet dataset, a derivation from the well-known and state-of-the-art NVBench dataset. The choice was made as this is the only dataset capable of supporting the generation and evaluation of the visualization, explanation (E), Captioning (C), and Suggestions for further data exploration (S). At the same time, additional datasets are made available, both in terms of chart collection methods discussed by Chen and Liu in their recent state-of-the-art report on corpora for automating chart generation [2] or the recently available VisText dataset for benchmarking captioning quality [48]. Although, as discussed in Section 5.2, these datasets are not readily usable for our purposes, we foresee the need to create an integrated resource capable of evaluating all the features produced by V-RECS in a combined way, possibly including also datasets like VIS30k [3] and OldOnlineVis [57].
User Evaluation: In its current form, the qualitative evaluation of V-RECS has been conducted through a comparative experiment with 15 expert users. We foresee the need to expand this study to a broader set of participants to explore the subtle differences that could exist in the presentation and significance layers of the EvaLLM evaluation stack and better evaluate the difference in our method between the proprietary GTP model and the open-source fine-tuned model. Given the richness of this activity, we foresee it as a dedicated effort to be conducted in future work.
Form of Explanation: In its current form, the explanation (E) provided to illustrate the generation process is reported only in textual form. As much as it is the standard for LLMs, and V-RECS aligns with it, we see as a limitation the lack of visual support to this explanation or lack of coordination between sub-parts of the explanation and visual elements of the generated chart. In this sense, the exploration of visual methods to support the explanation, as the ones reported by La Rosa et al. [21] or recently by Spinner et al. [43] for LLMs, should be the area for further enhancement of this approach.
7.2 Opportunities
Along with limitations, we also report opportunities that V-RECS enables for the visualization community:
Support to democratization of powerful LLMs The methodology followed for the creation of V-RECS can be reapplied to fine-tune a new model for different tasks in the visualization field, such as visualization modifications or generation of a bitmap version of a visualization, exploiting a large foundation model as the Teacher model (e.g., ChatGPT for text-to-text, or DALL-E for text -to-image) at a relatively low-cost. In this way, the capabilities of these larger models can be inherited (evaluating the degree of inheritance correctly) and used, analyzed, or modified by researchers and practitioners, fostering their spread and better comprehension.
Support to visualization and visual analytics pipelines V-RECS, the resulting model from the application of the proposed methodology, can be exploited for usage in visualization pipelines and visual analytics pipelines as a building block of the proposed system: looking at the former opportunity, tasks described by Schlesinger et al. [41] such as Chart recommendation and Rapid prototyping are already supported by V-RECS as is, while others such as the generation of Moodboards represent an easy extension. Interestingly, by construction V-RECS can also support tasks like visualizing the effect of training data on the outcomes of the model providing overall more controllability in the visualization recommendation process. Focusing on the latter, a very recent work by Zhao et al. [58], LEVA, systematized the usage of LLM for Visual Analytics through the definition of three main stages of intervention: onboarding, exploration, and summarization. In this context, V-RECS fits the summarization stage for generating visual summarization and explanations, eventually proposing directions to proceed with the analysis through its suggestions (S). It could even be adapted to the other two stages through the application of the methodology for generating a path of analysis (check the previous opportunity) by researchers or practitioners.
Support for training activities Still referring to the work by Zhao et al. [58], we foresee the opportunity to exploit V-RECS for training activities (referred to as the onboarding stage for Visual Analytics systems) on how to develop a visualization for a specific analysis task, through the support of explanations, or a workflow, exploiting a chain in which at each recommendation stage a new suggestion is selected and given as input for the next recommendation step. This activity would be in line with the recent trend of using LLM for education like the experiment by Chen et al. [6] testing ChatGPT-3.5 and 4 to complete assignments of a visualization course (CS171) and the preliminary work proposed by Joshi et al. [17] on leveraging ChatGPT-3.5 and Bard 2.0 for teaching a visualization technique.
Acknowledgements.
The authors wish to thank the participants who contributed to the manual assessment of V-RECS. This work was in part supported by a grant from Lazio Region, FESR Lazio 2021-2027, project @HOME (# F89J23001050007). This research has also been supported by the AWS Cloud Credit for Research Program, and the OpenAI’s Researcher Access Program.